Please answer with work

Find the area of the shaded region. 13) 13) y (0,4) y = 4 *२ 3 X (3,-5) A) 3 B) ၌ B) C) c) 33 23 3 D) 5
Answer #1

y=4-x² Soli- Area = (4-x²) du + (14-22) 4-42) du n 3 srea a 2 Area = (ax - 1 2 3 + 4x-2 (4627 - ) + (403) – 22)-(46) --9-6 Ar